OpenStudy (anonymous):

Umbuggio has two microphones which he uses to teach in RH100; one is Sony and one is Telefunken. The Sony fails with probability .3, while the Tele- funken fails with probability .4. On Monday morn- ing, Umbuggio picks a microphone at random. What is the probability that it will fail? (a) .7 (b) .35 (c) .12 (d) .58 (e) .42 (f) none of the above

geerky42 (geerky42):

 And you need to find \(P(pass')\), which is = \(P(pass'|Sony)P(Sony)+P(pass'|Telefunken)P(Telefunken)\\~\\=\left(.3\right)\left(\dfrac{1}{2}\right)+\left(.4\right)\left(\dfrac{1}{2}\right)\\=\cdots~?\)